Jack is driving at a speed that is 5 mph faster than Jill is driving. Jack drives 432 miles in the same amount of time that Jill drives 402 miles. How fast are Jack and Jill driving?
Answer by Alan3354(69443)

Jack is driving at a speed that is 5 mph faster than Jill is driving. Jack drives 432 miles in the same amount of time that Jill drives 402 miles. How fast are Jack and Jill driving?
------------------
They're 30 miles apart.
30/5 = 6 hours driving
-----------
432/6 = 72 mi/hr
402/6 = 67 mi/hr